Use the rules of differentiation to differentiate the following functions.



a. f(x)=2x²+6x


b.g(x)=7x⁴-3x²


c.y(x)=(4x)³- 18x²+6x


d.h(x)=(3x+4)²


e.h(x)=9x⅔+2/4√x

1
Expert's answer
2022-03-07T05:58:01-0500

a. "f(x)=2x^2+6x"

here we apply the sum rule

"d\/dx(f(x)) =d\/dx(g(x))+d\/dx(h(x))"

let "g(x)= 2x^2"

"h(x)=6x"

"=d\/dx(2x^3)+d\/dx(6x)"

"=6x^2+6"

b. g(x)= "7x^4-3x^2"

here we apply the difference rule

"d\/dx(g(x)) =d\/dx(f(x))-d\/dx(h(x))"

let "f(x)= 7x^4"

"h(x)=3x^2"

"=d\/dx(7x^4)-d\/dx(3x^2)"

"=28x^3-6x"

c. "y(x)=(4x)^3-18x^2+6x"

we start by simplifying the equation

= "(4x)^3=64x^3"

the simplified equation will be

"64x^3-18x^2+6x"

here we apply the sum/ difference rule

"d\/dx(y(x)) =d\/dx(f(x)) -d\/dx(g(x))+d\/dx(h(x))"

let "f(x)= 64x^3"

"g(x) = 18x^2"

"h(x)=6x"

"=d\/dx(y(x))=d\/dx(64x^3)-d\/dx(18x^2)+d\/dx(6x)"

"=192x^2-36x+6"

d. "h(x)=(3x+4)^2"

we apply the chain rule

"d\/dx[f(g(x)] =d\/d[g(x)][f(x)]*d\/dx(g(x))"

let f(x)= 2

"g(x)= 3x+4"

="2*(3x+4)*d\/dx(3x+4)"

= "2*(3x+4)*3"

="6(3x+4)"

= "18x+24"

e. "h(x)=9x"2/3+"2\/4\n\u200b\n \\sqrt{x}"

here we apply the sum rule.

d/dx(h(x))=d/dx(g(x))+d/dx(h(x))

="d\/dx(" 9x2/3)+"d\/dx" ("2\/4\n\u200b\n \\sqrt{x}")

= 6x -1/3-1/4x3/2